for y=2cos(4x-pi/3)
find
a)amplitude
b)period
c)phase shift

Answer by lwsshak3(11628)

for y=2cos(4x-pi/3)
find
a)amplitude
b)period
c)phase shift
**
Standard form of equation for cosine: y=Acos(Bx-C), A=amplitude, Period=2π/B,
Phase shift=C/B
For given problem:
a)amplitude: 2
b)period: 2π/B=2π/4=π/2
c)phase shift: C/B=(π/3)/4=π/12 (shift to the right)